OpenStudy (anonymous):

what is the product of x and y 12y + 4x= 180

OpenStudy (anonymous):

To find each variable, substitute 0 for each one. Start with finding y, so we add 0 in the x place. 12y + 4(0) = 180 12y + 0 = 180 Subtract 0 from both sides.. 12y = 180 Divide 12 by both sides. y = 15 Now we know y is 15, so substitute that for y and find x. 12(15) + 4x = 180 180 + 4x = 180 Subtract both sides by 180.. 4x = 0 Divide both sides by 4 to get: x = 0 Y = 15 X = 0 Does that help?